### v1

# Projection Theorems in the Presence of Expansions

Abstract.
We prove a restricted projection theorem for a certain one dimensional family of projections from $\mathbb{R}^{n}$ to $\mathbb{R}^{k}$.
The family we consider here arises naturally in the study of quantitative equidistribution problems in homogeneous dynamics.
1. Introduction
Restricted projection problems are intimately related to central questions in Fourier analysis and incidence geometry, and have been much studied, e.g., by Mattila, Falconer, Bourgain and others. More recently, certain restricted projection theorems have also found striking applications in homogeneous dynamics.
Let us recall the the classical Marstrand projection theorem: Let $K\subset\mathbb{R}^{n}$ be a compact subset, then for a.e. $v\in\mathbb{S}^{n-1}$
(1.1) $\dim{\rm p}_{v}(K)=\min(1,\dim K),$
where ${\rm p}_{v}(w)=w\cdot v$ is the orthogonal projection in the direction of $v$ and here and in what follows $\dim$ denotes the Hausdorff dimension. Analogous statements hold more generally for orthogonal projection into a.e. $m$-dimensional subspace, with respect to the Lebesgue measure on ${\rm Gr}(m,n)$.
Broadly speaking, restricted projection problems seek to obtain similar results as in (1.1) where $v$ is confined to a proper Borel subset $\mathsf{B}\subset\mathbb{S}^{n-1}$. Note, however, that without further restrictions on $\mathsf{B}$, (1.1) fails: e.g., if
$\mathsf{B}=\{(\cos t,\sin t,0):0\leq t\leq 2\pi\}$
is the great circle in $\mathbb{S}^{2}$ and $K$ is the $z$-axis, then ${\rm p}_{v}(K)=0$ for every $v\in\mathsf{B}$.
To avoid degenerations of this nature, one may, e.g., consider a curve $\gamma:[0,1]\to\mathbb{R}^{n}$ satisfying that
$\{\gamma^{\prime}(r),\gamma^{\prime\prime}(r),\ldots,\gamma^{(n)}(r)\}\qquad\text{spans $\mathbb{R}^{n}$ for all $r\in[0,1]$.}$
It is natural to inquire whether for almost all $r\in[0,1]$ the following holds
(1.2) $\dim{\rm p}_{\gamma^{\prime}(r)}(K)=\min(1,\dim K);$
again $K\subset\mathbb{R}^{n}$ is a compact subset.
Indeed (1.2) was conjectured by Fässler and Orponen [FO14] in dimension $3$. This conjecture was resolved by Käenmäki, Orponen, and Venieri [KOV17] and Pramanik, Yang, and Zahl [PYZ22] — these works rely on the work of Wolff and Schlag on circular Kakeya sets [Wol00, Sch03].
More recently, Gan, Guo, and Wang [GGW22] have established (1.2) in all dimensions using decoupling inequalities for the moment curve due to Bourgain, Demeter, and Guth [BDG16].
In this paper, we study a closely related family of projections. The consideration of these families are mainly motivated by the aforementioned applications to homogeneous dynamics.
Let us fix some notation in order to state the main results of this paper. For every $t\geq 0$ and $r\in[0,1]$, let $\pi_{t,r}:\mathbb{R}^{3}\to\mathbb{R}^{2}$ be
(1.3) $\pi_{t,r}(x,y,z)=\Bigl{(}e^{t}(x+ry+\tfrac{r^{2}}{2}z),y+rz\Bigr{)}$
The following is the main theorem in this paper.
1.1 Theorem.
Let $21\leq\alpha\leq 3/$ 2, and let $0<\delta_{0}\leq 1$. Let $F\subset B_{\mathbb{R}^{3}}(0,1)$ be a finite set satisfying the following:
#$\#(B_{\mathbb{R}^{3}}(w,\delta)\cap F)\leq C_{0}\cdot\delta^{\alpha}\cdot(\#F)\quad\text{for all $w\in\mathbb{R}^{3}$ and all $\delta\geq\delta_{0}$}$
where $C_{0}\geq 1$.
Let $1000<\varepsilon<\alpha/$ 100. For every $\delta\geq e^{t}\delta_{0}$, there is a subset $I_{\delta}\subset[1/2,1]$ with
$|[1/2,1]\setminus I_{\delta}|\ll_{\varepsilon}\delta^{\star\varepsilon}$
so that the following holds. Let $r\in I_{\delta}$, then there exists $F_{\delta,r}\subset F$ with
#$\#(F\setminus F_{\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon}\cdot(\#F)$
such that for all $w\in F_{\delta,r}$, we have
#($\#\Bigl{(}\{w^{\prime}\in F_{\delta,r}:\|\pi_{t,r}(w^{\prime})-\pi_{t,r}(w)|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}e^{-t/10}\delta^{\alpha-\varepsilon}\cdot(\#F)$
Remark.
Throughout the paper, the notation $a\ll b$ and $a^{\star b}$ mean $a\leq Db$ and $a^{Db}$, respectively, where $D$ is some positive constant whose dependence is explicated in different statements. Also, for a Borel subset $\mathsf{B}\subset\mathbb{R}^{d}$, we denote the Lebesgue measure of $\mathsf{B}$ by $|\mathsf{B}|$.
We also state a similar result in all dimensions $n\geq 3$, which relies on the aforementioned work of [GGW22].
Let $12\leq k\leq n-$ 1. For all $t\geq 0$ and all $r\in[0,1]$, let $\pi_{t,r}^{(k)}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ be
$\pi_{t,r}^{(k)}(w)=\Bigl{(}e^{t}w\cdot\xi^{\prime}(r),e^{t}w\cdot\xi^{\prime\prime}(r),\ldots,e^{t}w\cdot\xi^{(k-1)}(r),w\cdot\xi^{(k)}(r)\Bigr{)}$
where $w\cdot v$ is the usual inner product on $\mathbb{R}^{n}$ and
$\xi(r)=(\tfrac{r}{1!},\tfrac{r^{2}}{2!},\ldots,\tfrac{r^{n}}{n!})\subset\mathbb{R}^{n}.$
Using [GGW22, Thm. 2.1] and some elementary arguments, one can establish the following
1.2 Proposition.
Let $0<\alpha\leq k$, and let $0<\delta_{0}\leq 1$. Let $F\subset B_{\mathbb{R}^{n}}(0,1)$ be a finite set satisfying the following:
#$\#(B_{\mathbb{R}^{n}}(w,\delta)\cap F)\leq C_{0}\cdot\delta^{\alpha}\cdot(\#F)\quad\text{for all $w\in F$ and all $\delta\geq\delta_{0}$}$
where $C_{0}\geq 1$.
Let $1000<\varepsilon<\alpha/$ 100. For every $\delta\geq e^{t}\delta_{0}$, there is a subset $I_{\delta}\subset[1/2,1]$ with
$|[1/2,1]\setminus I_{\delta}|\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}$
so that the following holds. Let $r\in I_{\delta}$, then there exists $F_{\delta,r}\subset F$ with
#$\#(F\setminus F_{\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}\cdot(\#F)$
such that for all $w\in F_{\delta,r}$, we have
#($\#\Bigl{(}\{w^{\prime}\in F_{\delta,r}:\|\pi_{t,r}(w^{\prime})-\pi_{t,r}(w)|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}e^{-\alpha^{\prime}t}\delta^{\alpha-\varepsilon}\cdot(\#F)$
where $\alpha^{\prime}=\alpha$ if $10<\alpha\leq k-$ 1 and $\alpha^{\prime}=\alpha-1$ otherwise.
It is worth noting that Proposition 1.2 does not imply Theorem 1.1; indeed $\alpha^{\prime}=\alpha-1<1/10$ for $101\leq\alpha<11/$ 10. That said, it seems plausible that the approach developed in [GGW22] can be modified to prove a stronger version of Proposition 1.2 which then would imply (and perhaps improve) Theorem 1.1. However, our proof of Theorem 1.1 relies on much softer arguments and is an adaptation of [OV18].
Also as previously mentioned, our precise formulations here (which at least at first glance may appear somewhat ad hoc) are motivated by recent applications in homogeneous dynamics [LM23, LMW22, LMWY23].
2. Proof of Theorem 1.1
In this section we will prove Theorem 1.1. Let us begin by fixing some notation which will be used throughout this section.
Let $\mu$ denote the normalized counting measure on $F$:
$\mu(\mathsf{B})=\frac{\#(\mathsf{B}\cap F)}{\#F}\qquad\text{for any Borel set $\mathsf{B}\subset\mathbb{R}^{3}$}.$
Since $t$ is fixed throughout the argument, we will write $\pi_{r}$ for $\pi_{t,r}$. For every $w\in F$, all $r\in[0,1]$, and all $b>0$, let
$m^{b}(\pi_{r}(w))=\mu(\{w^{\prime}\in F:\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq b\}).$
More generally, given a subset $\mathsf{B}\subset\mathbb{R}^{3}$, let
$m^{b}(\pi_{r}(w)|\mathsf{B})=\mu(\{w^{\prime}\in\mathsf{B}:\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq b\}).$
For all $w\in F$ and all $b>0$, let $\mathsf{D}_{b}(w)=\{w^{\prime}:b\leq\|w-w^{\prime}\|\leq 2b\}$.
The following lemma is the main step in the proof of Theorem 1.1.
2.1 Lemma.
The following holds for all small enough $\eta$ and all large enough $C$. Let $F_{\rm hm}\subset F$ denote the set of $w\in F$ so that
|$|\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\}|\geq C\delta^{\eta}.$
Then $\mu(F_{\rm hm})\leq C\delta^{\eta}$.
We will use the following elementary lemma in the proof of Lemma 2.1.
2.2 Lemma.
Let $\sigma$ be a probability measure on $B(0,10)$ which satisfies
$\sigma(B(w,r))\leq\hat{C}r^{\beta}\quad\text{for all $w\in\mathbb{R}^{3}$ and all $r\geq r_{0}$}$
where $1/2\leq\beta\leq 1$ and $\hat{C}>0$.
The following holds for all $\varepsilon_{0}$ small enough. For every $b\geq r_{0}$, there exists a subset $E_{b}$ with $\sigma(B(0,10)\setminus E_{b})\leq b^{\varepsilon_{0}}$, and for every $z\in E_{b}$, there is a subset $I_{z}\subset[0,1]$ with $|[0,1]\setminus I_{z}|\leq b^{\varepsilon_{0}}$ so that for every $r\in I_{z}$
$\sigma\{z^{\prime}:|(1,r,\tfrac{r^{2}}{2})\cdot(z-z^{\prime})|\leq b\}\ll\hat{C}b^{49\beta/100}$
Proof.
This, rather weak estimate, follows from standard arguments, see e.g. [Ohm23, §3]. It is worth noting that much stronger result holds where $10049\beta/$ 100 is replaced by $\beta-o(1)$, see [LM23, Thm.  B.1] and [KOV17]. ∎
2.3 Lemma.
Let the notation and assumptions be as in Lemma 2.1. The following holds for all small enough $\eta$. There exists some $b\geq\delta^{1-3\eta}$ and a subset $F^{\prime}_{\rm hm}\subset F_{\rm hm}$ with
(2.1) $\mu(F^{\prime}_{\rm hm})\geq\delta^{3\eta/2}$
so that for all $w\in F^{\prime}_{\rm hm}$, we have
(2.2) |$|\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-7\eta}\}|\geq\delta^{3\eta/2}.$
Proof.
Let $b_{0}=\delta^{1-3\eta}$ and recall that $21\leq\alpha\leq 3/$ 2.
If $e^{-t}\geq\delta^{100\eta}$, then $e^{-t/10}\delta^{\alpha-18\eta}\geq\delta^{\alpha-8\eta}$. Therefore,
$\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\}\subset\\ \{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w))\geq C_{0}\delta^{\alpha-8\eta}\}.$
Now since $\mu(B(w,b_{0}))\leq C_{0}b_{0}^{\alpha}\leq C_{0}\delta^{\alpha-4.5\eta}$, there is some $b=b(w,r)\geq\delta^{1-3\eta}$ so that
(2.3) $m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-7.5\eta}.$
In view of this we assume $e^{-t}\leq\delta^{100\eta}$ for the rest of the argument. Put
$\Xi_{\delta}:=\{w:\mu(B(w,b_{0}))>C_{0}e^{-t/10}\delta^{\alpha-7\eta}\}$
Also for every $w\in F_{\rm hm}$, let
$I(w)=\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\cdot(\#F)\}.$
We now consider two possibilities:
Case 1: Assume $\mu(F_{\rm hm}\setminus\Xi_{\delta})\geq\delta^{\eta}/100$.
In this case, let $F^{\prime\prime}_{\rm hm}=F_{\rm hm}\setminus\Xi_{\delta}$. Then $\mu(F^{\prime\prime}_{\rm hm})\geq\delta^{\eta}/100$. Moreover, for every $w\in F^{\prime\prime}_{\rm hm}$, we have $\mu(B(w,b_{0}))\leq C_{0}e^{-t/10}\delta^{\alpha-7\eta}$.
Now since for every $r\in I(w)$, we have
$m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}$
there exists some $b=b(w,r)\geq\delta^{1-3\eta}$ so that
(2.4) $m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\gg C_{0}e^{-t/10}\delta^{\alpha-17\eta}.$
Case 2: Assume $\mu(F_{\rm hm}\setminus\Xi_{\delta})\leq\delta^{\eta}/100$.
In this case, we have $\mu(F_{\rm hm}\cap\Xi_{\delta})\geq\frac{C}{2}\delta^{\eta}$.
Fix a maximal $b_{0}/2$ separated subset $\{w_{1},\ldots,w_{M}\}$ of $F_{\rm hm}\cap\Xi_{\delta}$. Discarding a subset of $F_{\rm hm}$ with measure $\leq\mu(F_{\rm hm})/100$, we will assume that
$\mu(B(w_{i},b_{0}))\gg\mu(B(w_{i},4b_{0}))\qquad\text{ for all $1\leq i\leq M$},$
where the implied constant is absolute.
Let $\nu_{i}$ be the measure on $B(0,1)$ which is the image of
$\tfrac{1}{\mu(B(w_{i},b_{0}))}\mu|_{B(w_{i},b_{0})}$
under the map $w\to\frac{w_{i}-w}{b_{0}}$. Similarly, let $\tilde{\nu}_{i}$ denote the measure on $B(0,4)$ is the image of
$\tfrac{1}{\mu(B(w_{i},4b_{0}))}\mu|_{B(w_{i},4b_{0})}$
under the map $w\to\frac{w_{i}-w}{b_{0}}$.
Then for any $1\leq i\leq M$ and for $\sigma=\nu_{i},\tilde{\nu}_{i}$, we have
$\sigma(B(z,r))\leq e^{t/10}\delta^{-7\eta}r^{\alpha}\quad\text{for all $r\geq\delta_{0}/b_{0}$.}$
Thus, applying Lemma 2.2 with $\nu_{i},\tilde{\nu}_{i}$ (for any $1\leq i\leq M$), there exists a subset $E_{i}\subset B(w_{i},b_{0})$, with
$\nu_{i}(B(w_{i},b_{0})\setminus E_{i})\ll\delta^{\varepsilon_{0}}$
and for every $w\in E_{i}$, there is a subset $J_{w}$ with $|[0,1]\setminus J_{w}|\ll\delta^{\varepsilon_{0}}$ so that if $r\in J_{w}$, then for $\sigma=\frac{1}{\mu(B(w_{i},b_{0}))}\mu|_{B(w_{i},b_{0})}$ and $\sigma=\frac{1}{\mu(B(w_{i},4b_{0}))}\mu|_{B(w_{i},4b_{0})}$,
(2.5) $\sigma(\{w:|(1,r,\tfrac{r^{2}}{2})\cdot\tfrac{w-w^{\prime}}{b_{0}}|\leq e^{-t}\})$ $\ll e^{t/10}\delta^{-7\eta}e^{-49t/100}$
$\ll e^{-0.3t},$
where we used $e^{-t}\leq\delta^{100\eta}$ in the second inequality.
Let $F^{\prime\prime}_{\rm hm}=(F_{\rm hm}\cap\Xi_{\delta})\cap(\bigcup_{i}E_{i})$. Then
$\mu(F^{\prime\prime}_{\rm hm})\gg\delta^{\eta}$
For every $w\in F^{\prime\prime}_{\rm hm}$, let $\hat{I}(w)=I(w)\cap J_{w}$. Then $|\hat{I}(w)|\geq\frac{1}{2}\delta^{\eta}$. Moreover, for every $w\in F^{\prime\prime}_{\rm hm}$ and every $r\in\hat{I}(w)$,
$\{w^{\prime}\in B(w,b_{0}):\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq\delta\}\subset\\ \{w^{\prime}\in B(w_{i},4b_{0}):|(1,r,r^{2})\cdot\tfrac{w-w^{\prime}}{b_{0}}|\leq e^{-t}\}$
where $w\in B(w_{i},b_{0})$ (see (1.3) for the definition of $\pi_{r}=\pi_{t,r}$ and recall that $b_{0}=\delta^{1-3\eta}$). Applying (2.5) with $\sigma=\frac{1}{\mu(B(w_{i},4b_{0}))}\mu|_{B(w_{i},4b_{0})}$,
(2.6) $\mu(\{w^{\prime}\in B(w,b_{0}):\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq\delta\}\ll e^{-0.3t}\mu(B(w_{i},4b_{0}))$
$\ll C_{0}e^{-0.3t}b_{0}^{\alpha}<C_{0}e^{-0.2t}\delta^{\alpha-7\eta}$ ,
where in the last inequality, we used $e^{-t}\leq\delta^{100\eta}$.
In view of (2.6), for every $w\in F^{\prime\prime}_{\rm hm}$ and all $r\in\hat{I}(w)$, there exists some $b>b_{0}=\delta^{1-3\eta}$ so that (2.4) holds. That is:
(2.7) $m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\gg C_{0}e^{-t/10}\delta^{\alpha-17\eta}.$
Altogether, combining (2.3), (2.4), and (2.7), we have found a subset $F^{\prime\prime}_{\rm hm}\subset F_{\rm hm}$ with $\mu(F^{\prime\prime}_{\rm hm})\gg\delta^{\eta}$ and for every $w\in F^{\prime\prime}_{\rm hm}$ a subset $\hat{I}(w)\subset[\frac{1}{2},1]$ with $|\hat{I}(w)|\gg\delta^{\eta}$ so that the following holds. For every $w\in F^{\prime\prime}_{\rm hm}$ and all $w\in\hat{I}(w)$, there exists $b=b(w,r)\geq\delta^{1-3\eta}$ so that
$m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\gg C_{0}e^{-t/10}\delta^{\alpha-17\eta}$
Now applying pigeonhole principle (and Fubini’s theorem), there exists $b\geq\delta^{1-3\eta}$ and $F^{\prime}_{\rm hm}\subset F_{\rm hm}^{\prime\prime}$ with
$\mu(F^{\prime}_{\rm hm})\gg\delta^{3\eta/2}$
so that for all $w\in F^{\prime}_{\rm hm}$, we have
|$|\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-7\eta}\}|\geq\delta^{3\eta/2}.$
The proof is complete. ∎
Proof of Lemma 2.1.
Assuming $C$ is large enough, we may assume $\delta$ is small throughout the proof.
In view of Lemma 2.3, see in particular (2.1), we will replace $F_{\rm hm}$ by $F_{\rm hm}^{\prime}$ and assume that there is some $b\geq\delta^{1-3\eta}$ so that (2.2) holds for all $w\in F_{\rm hm}$.
For every $w\in F_{\rm hm}$, set
$I^{\prime}(w)=\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-7\eta}\}.$
Then $|I^{\prime}(w)|\gg\delta^{3/2}$, see (2.2). Choose three subsets $I^{\prime}_{j}(w)\subset\hat{I}^{\prime}(w)$ for $j=1,2,3$, which satisfy the following properties
(2.8) $|I^{\prime}_{j}(w)|\gg\delta^{2\eta}\qquad\text{and}\qquad{\rm dist}(I^{\prime}_{i}(w),I^{\prime}_{j}(w))\gg\delta^{3\eta/2}\quad\text{for $i\neq j$}$
For $j=1,2,3$, define
$E_{j}(w)=\{w^{\prime}\in F\cap\mathsf{D}_{b}(w):\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq\delta\text{ for some $r\in I^{\prime}_{j}(w)$}\}.$
We claim
(2.9) $\mu(E_{j}(w))\geq C_{0}be^{-t/10}\delta^{\alpha-1-6\eta}\qquad\text{for $j=1,2,3$.}$
Fix one $j$ and cover $I^{\prime}_{j}(w)$ with intervals $J_{1},\ldots,J_{N}$ of size $C^{\prime}\delta/b$ for some $C^{\prime}$ which will be chosen to be large. Thus
(2.10) $N\gg\delta^{2\eta}/(C^{\prime}\delta/b)=b\delta^{2\eta}/(C^{\prime}\delta)\gg b\delta^{2\eta-1}.$
Since $b\geq\delta^{1-3\eta}$, (2.10) in particular implies that $N\gg\delta^{-\eta}$.
For each $J_{i}$, let $r_{i}\in J_{i}\cap I^{\prime}_{j}(w)$. Discarding at most half of $J_{i}$’s, we will assume $|r_{i}-r_{i^{\prime}}|\geq C^{\prime}\delta/b$, and will continue to denote the collection by $J_{1},\ldots,J_{N}$. For every $1\leq i\leq N$, let
$E_{j,i}=\{w^{\prime}\in E_{j}(w):\|\pi_{r_{i}}(w)-\pi_{r_{i}}(w^{\prime})\|\leq\delta\}.$
Then $\mu(E_{j,i})\geq C_{0}e^{-\kappa t}\delta^{\alpha-7\eta}$. Moreover, $E_{j,i}\cap E_{j,i^{\prime}}=\emptyset$. Indeed if $w^{\prime}\in E_{j,i}\cap E_{j,i^{\prime}}$, then
$r_{i},r_{i^{\prime}}\in\{r\in[\tfrac{1}{2},1]:\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq\delta\}.$
Thus the set of $r\in[\tfrac{1}{2},1]$ so that
$|(1,r,\tfrac{r^{2}}{2})\cdot(w-w^{\prime})|\leq e^{-t}\delta\quad\text{and}$
$|(0,1,r)\cdot(w-w^{\prime})|\leq\delta$
has diameter $\geq C^{\prime}\delta/b$. Since $b\leq\|w-w^{\prime}\|\leq 2b$, we get a contradiction so long as $C^{\prime}$ is large enough.
Using (2.10) and $\mu(E_{j,i})\geq C_{0}e^{-\kappa t}\delta^{\alpha-7\eta}$, thus
$\mu(E_{j}(w))\gg N\cdot\mu(E_{j,i})\gg C_{0}be^{-t/10}\delta^{\alpha-1-5\eta},$
for $j=1,2,3$, as we claimed in (2.9).
Using $\mu(F_{\rm hm})\gg\delta^{3\eta/2}$ and (2.9), we conclude that
(2.11) $\mu(\{(w,w_{1},w_{2},w_{3})\in F_{\rm hm}\times F^{3}:w_{j}\in E_{j}(w)\})\gg\\ C_{0}^{3}b^{3}e^{-0.3t}\delta^{3\alpha-3-13\eta}$
We now find an upper bound for the measure of the set on the left side of (2.11). To that end, fix some $(w_{1},w_{2},w_{3})\in F^{3}$ so that there exists some $w\in F_{\rm hm}$ with $w_{j}\in E_{j}(w)$. In particular, we have $w_{j}\in\mathsf{D}_{b}(w)$, where $b\geq\delta^{1-3\eta}$, and
(2.12) $\|w_{1}-w_{j}\|\leq 4b,\qquad\text{for $j\in\{1,2,3\}$}.$
Let $r_{j}\in I^{\prime}_{j}(w)$ be so that $\|\pi_{r_{j}}(w)-\pi_{r_{j}}(w_{j})\|\leq\delta$. Then $|r_{i}-r_{j}|\geq\delta^{3\eta/2}$ for $i\neq j$ and
$(1,r_{j},r_{j}^{2})\cdot w=(1,r_{j},r_{j}^{2})\cdot w_{j}+O(e^{-t}\delta)\quad\text{for $j\in\{1,2,3\}$}.$
Thus $w$ belongs to a set with diameter $\ll e^{-t}\delta^{1-4.5\eta}$. This and (2.12) imply
$\mu(\{(w,w_{1},w_{2},w_{3})\in F_{\rm hm}\times F^{3}:w_{j}\in E_{j}(w)\})\ll C_{0}^{3}e^{-\alpha t}\delta^{\alpha-4.5\eta\alpha}b^{2\alpha}$
Comparing this upper bound and (2.11), we conclude
$b^{3}e^{-3t/10}\delta^{3\alpha-3-13\eta}\ll e^{-\alpha t}\delta^{\alpha-4.5\eta\alpha}b^{2\alpha},$
which implies
$\delta^{2\alpha-3-13\eta+4.5\eta\alpha}b^{3-2\alpha}\ll e^{(-\alpha+0.3)t}.$
Now using $b\geq\delta^{1-3\eta}$, $03-2\alpha>$ 0, and the above, we conclude that
(2.13) $\delta^{2\alpha-3-13\eta+4.5\eta\alpha}\delta^{(3-2\alpha)(1-3\eta)}=\delta^{-13\eta+4.5\eta\alpha-(9-6\alpha)\eta}\ll e^{(-\alpha+0.3)t}$
However, $-13\eta+4.5\eta\alpha-(9-6\alpha)\eta\leq-6\eta$, since $03-2\alpha>$ 0. Assuming $\delta$ is small enough and recalling that $3/10<\alpha$, (2.13) cannot hold. This contradiction completes the proof. ∎
Proof of Theorem 1.1
We now complete the proof of Theorem 1.1, which is based on Lemma 2.1 and Fubini’s theorem.
Recall that for every for all $r\in[\tfrac{1}{2},1]$ and all $w\in F$, we put
$m^{\delta}(\pi_{r}(w))=\mu\{w^{\prime}:\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq\delta\}$
Let $\varepsilon>0$ and let $\eta=\varepsilon/20$. For all $r\in[\frac{1}{2},1]$, let
$F_{\rm bad}(r)=\{w:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\}.$
The claim in Theorem 1.1 follows if we show that there is a subset $I_{\delta}\subset[\frac{1}{2},1]$ with $|[\frac{1}{2},1]\setminus I_{\delta}|\ll\delta^{\eta/2}$ so that for all $r\in I_{\delta}$, we have
$\mu(F_{\rm bad}(r))\ll\delta^{\eta/2}$
Let $C>1$, and assume that there exists a subset $I_{\rm bad}\subset[\frac{1}{2},1]$ with $|I_{\rm bad}|\geq C\delta^{\eta/2}$ so that for all $r\in I_{\rm bad}$, we have
$\mu(F_{\rm bad}(r))\geq C\delta^{\eta/2}.$
We will show this leads to a contradiction provided $C$ is large enough.
Equip $[\tfrac{1}{2},1]\times F$ with the product measure $\operatorname{Leb}\times\mu$. Let
$\mathsf{E}=\{(r,w)\in[\tfrac{1}{2},1]\times F:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\},$
and for every $w\in F$, let $\mathsf{E}_{w}=\{r:(r,w)\in\mathsf{E}\}$. The above then implies that
$\operatorname{Leb}\times\mu(\mathsf{E})\geq C^{2}\delta^{\eta}.$
Set $F^{\prime}=\{w\in F:|\mathsf{E}_{w}|\geq C\delta^{\eta}\}$. Then using Fubini’s theorem, we conclude
$\mu(F^{\prime})\geq\tfrac{1}{2}C^{2}\delta^{\eta}.$
Moreover, in in view of the definitions, for every $w\in F^{\prime}$
|$|\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\}|\geq C\delta^{\eta}$
This contradicts Lemma 2.1 provided $C$ is large enough.
The proof of complete. ∎
3. Proof of Proposition 1.2
Let us now turn to the proof of Proposition 1.2. As it was already mentioned, we will prove this theorem using [GGW22, Thm. 2.1]. Let us begin by recalling the notation used in  [GGW22, Thm. 2.1] and Proposition 1.2.
Recall that for all $r\in[0,1]$, we put $\xi(r)=(\tfrac{r}{1!},\tfrac{r^{2}}{2!},\ldots,\tfrac{r^{n}}{n!})\subset\mathbb{R}^{n}$. For all $1\leq k\leq n$, let $\mathfrak{p}_{r}^{(k)}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ be the orthogonal projection onto the space spanned by $\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}$. That is:
$\mathfrak{p}_{r}^{(k)}(w)=\Bigl{(}w\cdot\xi^{\prime}(r),w\cdot\xi^{(2)}(r),\ldots,w\cdot\xi^{(k)}(r)\Bigr{)}$
where $w\cdot v$ is the usual inner product on $\mathbb{R}^{n}$.
3.1 Theorem ([GGW22]).
Let $1\leq k\leq n$ and let $0<\alpha\leq k$. Let $\mu$ be the uniform measure on a finite set $F\subset B_{\mathbb{R}^{n}}(0,1)$ satisfying
$\mu(B_{\mathfrak{r}}(w,\delta)\cap\Theta)\leq C_{0}\delta^{\alpha}\qquad\text{for all $w$ and all $\delta\geq\delta_{0}$}$
where $C_{0}>0$.
Let $0<\varepsilon<10^{-4}\alpha$. For every $\delta\geq\delta_{0}$, there exists a subset $J_{\delta}\subset[0,1]$ with $|[0,1]\setminus J_{\delta}|\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}$ so that the following holds. Let $r\in J_{\delta}$, then there exists a subset $F_{\delta,r}\subset F$ with
$\mu(F\setminus F_{\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}$
such that for all $w\in F_{\delta,r}$ we have
$\mu\Bigl{(}\{w^{\prime}\in F:\|\mathfrak{p}_{r}^{(k)}(w)-\mathfrak{p}_{r}^{(k)}(w^{\prime})\|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}\delta^{\alpha-\varepsilon}$
Proof.
We deduce this from [GGW22, Thm. 2.1]. The argument is more or less standard. Indeed it is similar to the deduction of Theorem 1.1 from Lemma 2.1 and to (a finitary version of) the argument in [GGW22, §2].
Since $1\leq k\leq n$ is fixed throughout the argument, we will denote $\mathfrak{p}_{r}^{(k)}$ by $\mathfrak{p}_{r}$. Adapting the notation $m^{\delta}$ from the previous section: for every $r\in[\frac{1}{2},1]$ and all $w\in F$, put
$m^{\delta}(\mathfrak{p}_{r}(w))=\mu\{w^{\prime}:\|\mathfrak{p}_{r}(w)-\mathfrak{p}_{r}(w^{\prime})\|\leq\delta\}.$
Let $D_{1},\ldots$ be large constants which will be explicated later. Let $\eta=\varepsilon/D_{1}$. For all $r\in[\frac{1}{2},1]$, let
$F_{\rm bad}(r)=\{w:m^{\delta}(\mathfrak{p}_{r}(w))\geq C_{0}\delta^{\alpha-D_{1}\eta}\}.$
Assume contrary to the claim in Theorem 3.1 that there exists a subset $I_{\rm bad}\subset[\frac{1}{2},1]$ with $|I_{\rm bad}|\geq D_{2}\delta^{\eta^{2}/2}$ so that for all $r\in I_{\rm bad}$, we have
$\mu(F_{\rm bad}(r))\geq D_{2}\delta^{\eta^{2}/2}.$
We will get a contradiction with [GGW22, Thm. 2.1], provided that $D_{i}$’s are large enough.
First note that, for every $r\in I_{\rm bad}$, the number of $\delta$-boxes $\{\mathsf{B}_{i,r}\}$ required to cover $\mathfrak{p}_{r}(F_{\rm bad}(r))$ is $\leq D_{3}C_{0}^{-1}\delta^{-\alpha+D_{1}\eta}$. Following [GGW22], let $\mathcal{T}_{r}=\{\mathbb{T}_{i,r}\}$ where $\mathbb{T}_{i,r}=\mathfrak{p}_{r}^{-1}(\mathsf{B}_{i,r})\cap B_{\mathbb{R}^{n}}(0,1)$; note that
(3.1) #$\#\mathcal{T}_{r}\leq D_{3}C_{0}^{-1}\delta^{-\alpha+D_{1}\eta}.$
Select a maximal $\delta$-separated subset $\Lambda_{\delta}\subset I_{\rm bad}$ and extend this to a maximal $\delta$-separated subset $\hat{\Lambda}_{\delta}$ of $[\frac{1}{2},1]$. The definition of $F_{\rm bad}$
Let $\rho$ denote the uniform measure on $\hat{\Lambda}_{\delta}$, and as it was done in the proof of Theorem 1.1, equip $\hat{\Lambda}_{\delta}\times F$ with the product measure $\rho\times\mu$, and let
$\mathsf{E}$ =$=\{(r,w)\in\Lambda_{\delta}\times F:m^{\delta}(\pi_{r}(w))\geq C_{0}\delta^{\alpha-D_{1}\eta}\}$
=$=\{(r,w)\in\Lambda_{\delta}\times F:w\in F_{\rm bad}(r)\}.$
Then the above implies that $\rho\times\mu(\mathsf{E})\geq D_{2}^{2}\delta^{\eta^{2}}$.
For every $w\in F$, let $\mathsf{E}_{w}=\{r\in\Lambda_{\delta}:(r,w)\in\mathsf{E}\}$, and set
$F^{\prime}=\{w\in F:\rho(\mathsf{E}_{w})\geq D_{2}\delta^{\eta^{2}}\}.$
Then using Fubini’s theorem, we conclude $\mu(F^{\prime})\geq\tfrac{1}{2}D_{2}^{2}\delta^{\eta^{2}}$.
Recall that $\rho$ is the normalized counting measure on $\Lambda_{\delta}$. The above definitions thus imply
$\sum_{r\in\Lambda_{\delta}}1_{\mathbb{T}_{r}}(w)\geq D_{3}\delta^{\eta^{2}-1}\qquad\text{for all $w\in F^{\prime}$}$
where $D_{3}=O(D_{2})$ and the implied constant is absolute. Let $\mu^{\prime}$ denote the restriction of $\mu$ to $F^{\prime}$. Now by [GGW22, Thm 2.1], applied with $\eta^{2}$, $\mu^{\prime}$ and $\{\mathcal{T}_{r}:r\in\Lambda_{\delta}\}$, we have
$\sum_{r\in\Lambda_{\delta}}\#\mathcal{T}_{r}$ $\geq D_{4}(n,\varepsilon,\alpha)C_{0}^{-1}\mu^{\prime}(\mathbb{R}^{n})\delta^{-1-\alpha+D\eta}$
$\geq\tfrac{1}{2}D_{2}^{2}D_{4}(n,\varepsilon,\alpha)C_{0}^{-1}\delta^{\eta^{2}}\delta^{-1-\alpha+D\eta}$
where $D=10^{10n}$ and in the second line we used $\mu^{\prime}(\mathbb{R}^{n})=\mu(F^{\prime})\geq\tfrac{1}{2}D_{2}^{2}\delta^{\eta^{2}}$. Thus there exists some $r\in\Lambda_{\delta}$ so that
(3.2) #$\#\mathcal{T}_{r}\geq\tfrac{1}{2}C_{0}^{-1}D_{4}(n,\varepsilon,\alpha)D_{2}^{2}\delta^{-\alpha+(D+1)\eta}.$
Now comparing (3.2) and (3.1) we get a contradiction so long as $D_{1}$ is large enough and $\delta$ is small enough. The proof is complete. ∎
We now turn to the proof of Proposition 1.2.
Proof of Proposition 1.2.
Let $\delta$ and $k$ be as in Proposition 1.2. As before, we let $\mu$ be the uniform measure on a finite set $F$.
Apply Theorem 3.1 with $\delta^{\prime}:=e^{-t}\delta\geq\delta_{0}$ and with $k$ and $k-1$. Then for every $r\in J_{\delta}$ and $w\in F_{\delta^{\prime},r}$, we have
(3.3) $\mu\Bigl{(}\{w^{\prime}\in F:\|\mathfrak{p}_{r}^{(\ell)}(w)-\mathfrak{p}_{r}^{(\ell)}(w^{\prime})\|\leq e^{-t}\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}(e^{-t}\delta)^{\min(\alpha,\ell)-\varepsilon}$
for $\ell=k,k-1$.
Also note that if $10<\alpha\leq k-$ 1, then
$\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\}\subset\\ \{w^{\prime}\in F:\|\mathfrak{p}_{r}^{(k-1)}(w)-\mathfrak{p}_{r}^{(k-1)}(w^{\prime})\|\leq e^{-t}\delta\}.$
Thus (3.3), with $\ell=k-1$ implies
(3.4) $\mu(\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\})\ll_{\varepsilon}C_{0}e^{-\alpha t}\delta^{\alpha-\varepsilon}$
as we claimed.
Assume now that $k-1<\alpha\leq k$. Let $Q\subset{\rm Span}\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}\simeq\mathbb{R}^{k}$ denote the box $\delta^{\prime}\times\cdots\times\delta^{\prime}\times\delta$ centered at the origin, where directions correspond directions of $\xi^{(i)}(r)$. That is: $Q$ is box of size $\delta^{\prime}$ in the direction of $\xi^{(i)}(r)$ for all $11\leq i\leq k-$ 1 and of size $\delta$ in the direction of $\xi^{(k)}(r)$. We cover
$Q^{\prime}:=(Q+\pi_{t,r}(w))\cap\pi_{t,r}(F_{\delta^{\prime},r})$
with $\ll e^{t}$ many boxes $B_{i}$ of size $\delta^{\prime}$. Then (3.3), applied with $2\delta^{\prime}$, implies
(3.5) $\mu(Q^{\prime})$ $\ll_{\varepsilon}(C_{0}(2\delta^{\prime})^{\alpha-\varepsilon})\cdot e^{t}$
$\ll_{\varepsilon}C_{0}e^{(1-\alpha)t}\delta^{\alpha-\varepsilon}.$
Finally, note that
$\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\}\subset Q+\pi_{t,r}(w).$
This and (3.5) complete the proof in this case as well. ∎